OpenStudy (anonymous):

8^x/8^2=8^-2 Use the properties of exponents to simplify the left side of the equation. Then solve the equation.

OpenStudy (experimentx):

8^x/8^2=8^-2 => 8^(x-2) = 8^-2 => x-2 = -2 or x=0